Shift Summary for Evening shift, Nov 2, 1999

Miguel Barrio (mbarrio@hep.uchicago.edu)
Wed, 3 Nov 1999 00:01:31 -0600

Shift Summary for Evening shift, Nov 2, 1999

On Shift: Rick Kessler, Miguel Barrio

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of 99 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:

No beam this shift.

Activities:

circa 17:00 Had problems with the HCC, Sydney and BobT worked on it for
a while. After a while of futzing with it they managed to
get back to the state things were on this morning.

21:00 HCC problems seem to have cleared. The event display shows
that the upper-right quadrant of the CsI has dropped out.
Doing an init cleared the problem.

21:30 BobT is going to Burger King to get some food. He looks at
me with amazement and pitty when I ask him to get me a "BK
Broiler with NO MAYO." Oh well...

22:00 Permit gets lost somewhere. Followed it all the way from
crate 53 to the fruits. Somehow, futzing with the connectors
seems to have fixed the problem.

=========================================================================
Runs ended this shift: none

Tapes brought to FCC this shift: 0